OpenStudy (anonymous):

3 2 - 9 A. 8 --- 729 B. 6 - 9 C. 2 --- 729 D. 8 - 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If it is \(\huge (\frac{2}{9})^3\), then evaluate as \(\huge \frac{2^3}{9^3}\),

OpenStudy (anonymous):

so 8 over 9

OpenStudy (anonymous):

2^3 = 8 9^3= ?????

OpenStudy (anonymous):

729

OpenStudy (anonymous):

So it's... ?

OpenStudy (anonymous):

so 8 over 729

OpenStudy (anonymous):

Good!